i dont know how to solve this can you help the problem is -2(x-6)+7=35 solve the equasion and check your solution and i have to find out what x is
Answer by xcentaur(357)

-2(x-6)+7=35
-2x+12+7=35
-2x=35-12-7
-2x=35-19
x=16/-2
x=-8


Prabhat